OpenStudy (anonymous):

A random number generator draws at random with replacement from the digits 0, 1, 2, 3, 4, 5, 6, 7, 8, 9. Find the chance that the digit 5 appears on more than 11% of the draws, if 1. 100 draws are made 2. 1000 draws are made

OpenStudy (perl):

ok think of this as a binomial problem. you are drawing 100 numbers randomly , the success if picking 5 , a failure is picking a different number. probability of success (picking a 5) is 1/10. probability of failure is 9/10 so we have a binomial distribution. now i think you can approximate it with a normal distribution

OpenStudy (perl):

the mean of a binomial distribution is mean = n*p std dev = sqrt ( n p q )

OpenStudy (perl):

now you want probability ( X > = 11% of the 100 draws) = P ( X > = 11)

OpenStudy (perl):

P( X > = 11) = 1 - P ( X < 11) =.4168

OpenStudy (perl):

and you can solve this using a normal distribution approximation to the binomial distribution the normal distribution will have mean = n * p , and st dev = sqrt(n*p*q)

OpenStudy (anonymous):

Unfortunately, I haven't got correct results.

OpenStudy (anonymous):

0.312

OpenStudy (anonymous):

i THINK , ANSWER 1

OpenStudy (anonymous):

sorry , I dont understand

OpenStudy (anonymous):

pls any one can ans above askd question clearly